OpenStudy (anonymous):

write the following square root in the form a√b= where a and b are integers and b has the least value possible.√63=

OpenStudy (anonymous):

\[\LARGE \sqrt{63}=\sqrt{9\cdot 7}=\sqrt9 \cdot \sqrt7=3\sqrt7\]

OpenStudy (anonymous):

Thanks Kreshnik. Have one more question for you to help me with if you can.

OpenStudy (anonymous):

Find the prime factorization of 55^5*65*9^15

OpenStudy (anonymous):

why don't you post it on wall... what if I don't know it ?

OpenStudy (anonymous):

auff...

OpenStudy (anonymous):

prime of 120 does it mean: 60*2 30*2*2 15*2*2 ... ??

OpenStudy (anonymous):

Ok thnaks

OpenStudy (anonymous):

no no .. wait, that's not the answer :F lol I was just asking what does prime mean. Because I'm not american I don't understand well enough sometimes :( @dog1985

OpenStudy (anonymous):

Ok. Hold on I have the defintion for prime factor.

OpenStudy (anonymous):

@dog1985 I hope it's not what I think it is.. do you know that 55^5*65*9^15 is going to be a very LOOOONG... number ! ? :O

OpenStudy (anonymous):

It gave me 4 to choose from. A. 3^15*5^6*11^5*13 B. 3^30*5^6*11^5*13 C. 3^30*5^6*11*13 D. 3^30*5^5*11^5*13

OpenStudy (anonymous):

\[\LARGE 9^{15}=\left(3^2\right)^{15}=3^{30}\] so let's eliminate option A. :) ... now let me think ;)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

\[\LARGE 55^{5}=(5\cdot 11)^5=5^5\cdot 11^5\] so we have.. \[\LARGE 55^5\cdot 65\cdot 9^{15}\] \[\LARGE (5\cdot 11)^5\cdot 65\cdot 9^{15}\] \[\LARGE 5^5\cdot 11^5\cdot 65\cdot 3^{30}\] now we have.. \[\LARGE 65=5\cdot 13\] so.. \[\LARGE 5^5\cdot 11^5\cdot (5\cdot 13)\cdot 3^{30}\] \[\LARGE 5^{5+1 }\cdot 11^5\cdot 13\cdot 3^{30}\] \[\LARGE 5^{6 }\cdot 11^5\cdot 13\cdot 3^{30}\] so option B looks to me correct since it is.. 3^30*5^6*11^5*13 with latex it looks.. \[\LARGE 3^{30}*5^6*11^5*13\] SO i GUESS THEY ARE THE SAME ;) OPTION B DEFINITELY !

OpenStudy (anonymous):

You helped me understand it better than I did before. Thanks for your help.

OpenStudy (anonymous):

My pleasure :) ... You're welcome ;)